# Algebraic Barriers to Halving Algorithmic Information Quantities in Correlated Strings

### v1

@@ 1.1 The main construction
(where $\langle x,y\rangle$ is the code of the pair $(x,y)$), and prove that for every $z$ satisfying $C(a|z)\approx\frac{1}{2}C(a),\ C(b|z)\approx\frac{1}{2}C(b),$ the value of $C(c|z)$ must be much bigger than $\frac{1}{2}C(c)$, see Theorem 5.
@@ 3.1 Typical pairs: interface between the combinatorial language and AIT
$\left\{\begin{array}[]{l}C(x)\stackrel{{\scriptstyle\log}}{{=}}\alpha+\gamma+O(\log n),\\ C(y)\stackrel{{\scriptstyle\log}}{{=}}\beta+\gamma+O(\log n),\\ C(x,y)\stackrel{{\scriptstyle\log}}{{=}}\alpha+\beta+\gamma+O(\log n).\end{array}\right.$ (6)
@@
$|E^{\prime}|\geq 2^{C(xy|z)-O(\log n)}=2^{\alpha+\beta+\gamma-O(\log n)},$
@@ 3.2 Typical incidences in a projective plane
$C(x)=2n+O(\log n),\ C(y)=2n+O(\log n),\ C(x,y)=3n+O(\log n),$
@@
$C(x|z)=n+O(\log n),\ C(y|z)=n+O(\log n),\ C(x,y|z)=1.5n+O(\log n)$
@@
$C(x|y,z)=0.5n+O(\log n),\ C(y|x,z)=0.5n+O(\log n),\ I(x:y|z)=0.5n+O(\log n),$
@@
$C(x)=2n+O(\log n),\ C(y)=2n+O(\log n),\ C(x,y)=3n+O(\log n),$
@@
$C(x|y,z)=(0.5-\epsilon)n+O(\log n),\ C(y|x,z)=(0.5-\epsilon)n+O(\log n)$ (7)
we have $I(x:y|z)\geq\frac{4}{7}n-O(\epsilon n)-O(\log n)\gg 0.5n$.
@@
$|L^{\prime}|^{7/8}<|R^{\prime}|<|L^{\prime}|^{8/7}\text{ and }\max\{|L^{\prime}|,|R^{\prime}|\}\leq\mathbb{F}^{8/7}$
@@
$\begin{array}[]{rcl}C(a|z)&=&\frac{1}{2}C(a)+O(\log n),\\ C(b|z)&=&\frac{1}{2}C(b)+O(\log n),\\ C(c|z)&=&\frac{1}{2}C(c)+O(\log n).\end{array}$
@@ 4.2 Prime field: a negative result
    $C(w|\langle m_{A},m_{B}\rangle)\geq k-O(\log n),$
@@
Then $k\leq 3n/7+O(\epsilon n)+O(\log n).$ Thus, for small enough $\epsilon$, the size of the secret key $k\ll n/2$.
@@
$C(x|t,y)=(0.5-\epsilon)n+O(\log n)\text{ and }C(y|t,x)=(0.5+\epsilon)n+O(\log n).$
@@
$I(x:y|z)\stackrel{{\scriptstyle\log}}{{=}}I(x:y|t)-C(w)+O(\log n)\stackrel{{\scriptstyle\log}}{{\leq}}n-k,$
@@ 4.3 Field with a large subfield: a positive result
    $C(w|\langle m_{A},m_{B}\rangle)\geq n/2-O(\log n),$
@@
Since the incidence $(x,y)$ is typical , i.e., $C(x,y)\stackrel{{\scriptstyle\log}}{{=}}3n$, we obtain $C(f|m_{1},m_{2})\stackrel{{\scriptstyle\log}}{{\geq}}n/2$. ∎
@@ A.1 Halving complexities with a hashing argument
  • $C(r)=k$,
  • $C(r|x)=O(\log n)$ and
  • $I(r:y)=\max\{0,k-C(x|y)\}+O(\log n)$.
@@
First of all, we apply Proposition 3 and obtain a string $r^{\prime}$ of length $C(x|y)$ such that $C(r^{\prime}|x)=O(\log n)$ and $C(x|r^{\prime},y)=O(\log n)$. The latter condition means that the difference between $C(x|y)$ and $C(x|r^{\prime},y)$ is equal to $C(x|y)=C(r^{\prime})$. Therefore,
$I(r^{\prime}:y)=O(\log n).$
If $k\leq C(x|y)$, we define $r$ as the prefix of $r^{\prime}$ of length $k$, and we are done. Otherwise, we apply again Proposition 3 and obtain $r^{\prime\prime}$ of length $C(x)-C(x|y)$ such that $C(r^{\prime\prime}|x)=O(\log n)$ and $C(x|r^{\prime},r^{\prime\prime})=O(\log n)$. Let $r^{\prime\prime\prime}$ be the prefix of $r^{\prime\prime}$ of length $k-C(x|y)$. We set $r=r^{\prime}r^{\prime\prime\prime}$ (the concatenation of two strings). It is easy to see that
$C(r)=C(r)+C(r^{\prime})+O(\log n)=k+O(\log n)$
@@
$C(r|y)=C(r^{\prime}|y)+O(\log n)=C(x|y)+O(\log n),$
which implies $I(r:y)=k-C(x|y)+O(\log n)$. ∎
@@
$C(a)=\alpha+\gamma+O(\log n),\ C(b)=\beta+\gamma+O(\log n).$
@@
  • $C(p|a)=O(\log n)$ (and, therefore, $I(p:a)=C(p)=\frac{\alpha+\gamma}{2}$),
  • $C(p|b)=C(p)+O(\log n)$ (in other words, $I(p:b)=O(\log n)$).
@@
  • $C(q)=\frac{\beta+\gamma}{2}$,
  • $C(q|b)=O(\log n)$ (and, therefore, $I(q:b)=C(q)=\frac{\beta+\gamma}{2}$),
  • $C(q|a)=C(q)+O(\log n)$ (thus, $I(q:p)\leq I(q:a)+O(\log n)=O(\log n)$).
We let $z=pq$ (concatenation of two strings). A straightforward computation gives
$\begin{array}[]{rcl}C(a|z)&=&C(a|pq)=C(a,p,q)-C(p,q)\\ &=&C(a,q)-C(p,q)=C(a)+C(q)-C(p)-C(q)\\ &=&C(a)-C(p)=\alpha+\gamma-\frac{\alpha+\gamma}{2}=\frac{\alpha+\gamma}{2}=\frac{1}{2}C(a).\end{array}$
$\begin{array}[]{rcl}C(b|z)&=&C(b|pq)=C(b,p,q)-C(p,q)\\ &=&C(b,p)-C(p,q)=C(b)+C(p)-C(p)-C(q)\\ &=&C(b)-C(q)=\beta+\gamma-\frac{\beta+\gamma}{2}=\frac{\beta+\gamma}{2}=\frac{1}{2}C(b).\end{array}$
@@
$I(p:y)=C(p)-C(a|b)=\frac{\alpha+\gamma}{2}-\alpha=\frac{\gamma-\alpha}{2}.$
@@
  • $C(q)=\frac{\alpha+\beta}{2}$,
  • $C(q|b)=O(\log n)$ (and, therefore, $I(q:b)=C(q)=\frac{\alpha+\beta}{2}$),
  • $C(q|a)=C(q)+O(\log n)$ (thus, $I(q:p)\leq I(q:a)+O(\log n)=O(\log n)$).
Similarly to Case 1 we let $z=pq$ and obtain the necessary conditional complexities: the computation
$\begin{array}[]{rcl}C(a|z)=\frac{\alpha+\gamma}{2}=\frac{1}{2}C(a).\end{array}$
@@
$\begin{array}[]{rcl}C(b|z)&=&C(b|pq)=C(b,p,q)-C(p,q)\\ &=&C(b,p)-C(p,q)=C(b|p)+C(p)-C(p)-C(q)\\ &=&C(b)-I(p:b)-C(q)=\beta+\gamma-\frac{\gamma-\alpha}{2}-\frac{\alpha+\beta}{2}=\frac{\beta+\gamma}{2}=\frac{1}{2}C(b).\end{array}$
@@ A.2 Halving complexities with a topological argument
Theorem 8.
@@
Then there exists a string $z$ such that
$C(x|z)=n+O(\log n),\ C(y|z)=n+O(\log n).$
@@
We will show that the required $z$ can be constructed as a pair $a=\langle x^{\prime},y^{\prime}\rangle$, where $x^{\prime}$ and $y^{\prime}$ are prefixes of $x$ and $y$ respectively. We are going to prove that such prefixes of $x$ and $y$ can be chosen properly, although we cannot specify their lengths explicitly
@@
First of, we observe that the points $(0,0)$, $(|x|,0)$, $(0,|y|)$, $(|x|,|y|)$ are mapped to the ($O(\log n)$-neighborhoods of) points
@@
on the right [^3] For example, $(\alpha,\beta)=(|x|,0)$ means that $z$ is a pair consisting of the entire $x$ and the empty prefix of $y$; conditional on this $z$, complexity of $x$ vanishes, and conditional complexity of $y$ shrinks to $C(y|x)=n\pm O(\log n)$. The other three cases can be settled in a similar way..
When we go along the vertical line $(|x|,0)$$(|x|,|y|)$, the string $z$ consists of the entire $x$ and a gradually growing prefix of $y$. Accordingly, the value of $C(x|z)$ remains equal to $O(1)$, and $C(y|z)$ decreases from $C(y|z)\pm O(\log n)$ to 00. That is, the image of $(\alpha,\beta)$ (taken from the left picture) runs through the segment $(0,n)$$(0,0)$ (on the right picture). Similarly, when $(\alpha,\beta)$ run the segment $(0,|y|)$$(|x|,|y|)$ on the left, the images run through the segment $(n,0)$$(0,0)$ on the right, at least within $O(\log n)$ precision.
Now let $(\alpha,\beta)$ runs through the segment $(0,0)$$(|x|,0)$. We know that the image of $(\alpha,\beta)$ (i.e., the corresponding values of $(C(x|z),C(y|z))$) gradually changes from $(2n,2n)$ to $(0,n)$. However, we cannot control exactly the behavior of $(C(x|z),C(y|z))$ on this interval: the image of $(0,0)$$(|x|,0)$ does not need to be a straight line; we only know that it is slowly changing line that descends from $(2n,2n)$ to $(0,n)$ and remains all the way above the horizontal line $C(y|z)=n-O(\log n)$ (indeed, the complexity value $C(y|x^{\prime})$ cannot be be much smaller than $C(y|x)=n-O(\log n)$).
Similarly, when $(\alpha,\beta)$ runs through the segment $(0,0)$$(0,|y|)$, the pair of values $(C(x|z),C(y|z))$ goes from $(2n,2n)$ to $(n,0)$. Thus, the border line of the rectangle on the left in Fig. 3 is mapped to the bold line on the right in the same figure. Observe that the point $(n,n)$ is inside (at least within $O(\log n)$ precision) the loop shown in bold on the right in Fig. 3. It remains to show that some point from the domain of the mapping (13) is mapped to the $O(1)$-neighborhood of $(n,n)$. In what follows we do it in three steps.

### v2

@@ 1.1 The main construction
(where $\langle x,y\rangle$ is the code of the pair $(x,y)$), and prove that for every string $z$ satisfying the conditions $C(a|z)\approx\frac{1}{2}C(a)\text{ that }C(b|z)\approx\frac{1}{2}C(b),$ the value of $C(c|z)$ must be much bigger than $\frac{1}{2}C(c)$, see Theorem 5.
@@ 3.1 Typical pairs: interface between the combinatorial language and AIT
$\left\{\begin{array}[]{l}C(x)\stackrel{{\scriptstyle\log}}{{=}}\alpha+\gamma,\\ C(y)\stackrel{{\scriptstyle\log}}{{=}}\beta+\gamma,\\ C(x,y)\stackrel{{\scriptstyle\log}}{{=}}\alpha+\beta+\gamma.\end{array}\right.$ (6)
@@
$|E^{\prime}|\geq 2^{C(x,y|z)-O(\log n)}=2^{\alpha+\beta+\gamma-O(\log n)},$
@@ 3.2 Typical incidences in a projective plane
$C(x)\stackrel{{\scriptstyle\log}}{{=}}2n,\ C(y)\stackrel{{\scriptstyle\log}}{{=}}2n,\ C(x,y)\stackrel{{\scriptstyle\log}}{{=}}3n,$
@@
$C(x|z)\stackrel{{\scriptstyle\log}}{{=}}n,\ C(y|z)\stackrel{{\scriptstyle\log}}{{=}}n,\ C(x,y|z)\stackrel{{\scriptstyle\log}}{{=}}1.5n$
@@
$C(x|y,z)\stackrel{{\scriptstyle\log}}{{=}}0.5n,\ C(y|x,z)\stackrel{{\scriptstyle\log}}{{=}}0.5n,\ I(x:y|z)\stackrel{{\scriptstyle\log}}{{=}}0.5n,$
@@
$C(x)\stackrel{{\scriptstyle\log}}{{=}}2n,\ C(y)\stackrel{{\scriptstyle\log}}{{=}}2n,\ C(x,y)\stackrel{{\scriptstyle\log}}{{=}}3n,$
@@
$C(x|y,z)\stackrel{{\scriptstyle\log}}{{=}}(0.5-\epsilon)n,\ C(y|x,z)\stackrel{{\scriptstyle\log}}{{=}}(0.5-\epsilon)n$ (7)
we have $I(x:y|z)\stackrel{{\scriptstyle\log}}{{\geq}}\frac{4}{7}n-O(\epsilon n)\gg 0.5n$.
@@
$|L^{\prime}|^{7/8}<|R^{\prime}|<|L^{\prime}|^{8/7}\text{ and }\max\{|L^{\prime}|,|R^{\prime}|\}\leq|\mathbb{F}|^{8/7}$
@@
$\begin{array}[]{rcl}C(a|z)&\stackrel{{\scriptstyle\log}}{{=}}&\frac{1}{2}C(a),\\ C(b|z)&\stackrel{{\scriptstyle\log}}{{=}}&\frac{1}{2}C(b),\\ C(c|z)&\stackrel{{\scriptstyle\log}}{{=}}&\frac{1}{2}C(c).\end{array}$
@@ 4.2 Prime field: a negative result
    $C(w|\langle m_{A},m_{B}\rangle)\stackrel{{\scriptstyle\log}}{{\geq}}k,$
@@
Then $k\stackrel{{\scriptstyle\log}}{{\leq}}3n/7+O(\epsilon n).$ Thus, for small enough $\epsilon$, the size of the secret key $k\ll n/2$.
@@
$C(x|t,y)\stackrel{{\scriptstyle\log}}{{=}}(0.5-\epsilon)n\text{ and }C(y|t,x)\stackrel{{\scriptstyle\log}}{{=}}(0.5+\epsilon)n.$
@@
$I(x:y|z)\stackrel{{\scriptstyle\log}}{{=}}I(x:y|t)-C(w)\stackrel{{\scriptstyle\log}}{{\leq}}n-k,$
@@ 4.3 Field with a large subfield: a positive result
    $C(w|\langle m_{A},m_{B}\rangle)\stackrel{{\scriptstyle\log}}{{\geq}}n/2,$
@@
Since the incidence $(x,y)$ is typical , i.e., $C(x,y)\stackrel{{\scriptstyle\log}}{{=}}3n$, we obtain $C(f|m_{1},m_{2})\stackrel{{\scriptstyle\log}}{{\geq}}n/2$. Thus, $C(f|m_{1},m_{2})\stackrel{{\scriptstyle\log}}{{=}}C(f)$, and the claim is proven. ∎
@@ A.1 Halving complexities with a hashing argument
  • $C(r)\stackrel{{\scriptstyle\log}}{{=}}k$,
  • $C(r|x)\stackrel{{\scriptstyle\log}}{{=}}0$, and
  • $I(r:y)\stackrel{{\scriptstyle\log}}{{=}}\max\{0,k-C(x|y)\}$.
@@
First of all, we apply Proposition 3 and obtain a string $r^{\prime}$ of length $C(x|y)$ such that $C(r^{\prime}|x)\stackrel{{\scriptstyle\log}}{{=}}0$ and $C(x|r^{\prime},y)\stackrel{{\scriptstyle\log}}{{=}}0$. The latter condition means that the difference between $C(x|y)$ and $C(x|r^{\prime},y)$ is equal to $C(x|y)=C(r^{\prime})$. Therefore, $I(r^{\prime}:y)\stackrel{{\scriptstyle\log}}{{=}}0.$ If $k\leq C(x|y)$, we define $r$ as the prefix of $r^{\prime}$ of length $k$, and we are done. Otherwise, we apply again Proposition 3 and obtain $r^{\prime\prime}$ of length $C(x)-C(x|y)$ such that $C(r^{\prime\prime}|x)\stackrel{{\scriptstyle\log}}{{=}}0$ and $C(x|r^{\prime},r^{\prime\prime})\stackrel{{\scriptstyle\log}}{{=}}0$. Let $r^{\prime\prime\prime}$ be the prefix of $r^{\prime\prime}$ of length $k-C(x|y)$. We set $r=\langle r^{\prime},r^{\prime\prime\prime}\rangle$ (encoding of the pair of strings). It is easy to see that
$C(r)\stackrel{{\scriptstyle\log}}{{=}}C(r)+C(r^{\prime})\stackrel{{\scriptstyle\log}}{{=}}k$
@@
$C(r|y)\stackrel{{\scriptstyle\log}}{{=}}C(r^{\prime}|y)\stackrel{{\scriptstyle\log}}{{=}}C(x|y),$
which implies $I(r:y)\stackrel{{\scriptstyle\log}}{{=}}k-C(x|y)$. ∎
@@
$C(a)\stackrel{{\scriptstyle\log}}{{=}}\alpha+\gamma,\ C(b)\stackrel{{\scriptstyle\log}}{{=}}\beta+\gamma.$
@@
  • $C(p|a)\stackrel{{\scriptstyle\log}}{{=}}0$ (and, therefore, $I(p:a)=C(p)=\frac{\alpha+\gamma}{2}$),
  • $C(p|b)\stackrel{{\scriptstyle\log}}{{=}}C(p)$ (in other words, $I(p:b)\stackrel{{\scriptstyle\log}}{{=}}0$).
@@
  • $C(q)\stackrel{{\scriptstyle\log}}{{=}}\frac{\beta+\gamma}{2}$,
  • $C(q|b)\stackrel{{\scriptstyle\log}}{{=}}O(\log n)$ (and, therefore, $I(q:b)\stackrel{{\scriptstyle\log}}{{=}}C(q)\stackrel{{\scriptstyle\log}}{{=}}\frac{\beta+\gamma}{2}$),
  • $C(q|a)\stackrel{{\scriptstyle\log}}{{=}}C(q)$ (thus, $I(q:p)\stackrel{{\scriptstyle\log}}{{\leq}}I(q:a)\stackrel{{\scriptstyle\log}}{{=}}0$).
We let $z=\langle p,q\rangle$ (encoding of the pair ofstrings). A straightforward computation gives
$\begin{array}[]{rcl}C(a|z)&\stackrel{{\scriptstyle\log}}{{=}}&C(a|pq)\stackrel{{\scriptstyle\log}}{{=}}C(a,p,q)-C(p,q)\\ &\stackrel{{\scriptstyle\log}}{{=}}&C(a,q)-C(p,q)\stackrel{{\scriptstyle\log}}{{=}}C(a)+C(q)-C(p)-C(q)\\ &\stackrel{{\scriptstyle\log}}{{=}}&C(a)-C(p)\stackrel{{\scriptstyle\log}}{{=}}\alpha+\gamma-\frac{\alpha+\gamma}{2}\stackrel{{\scriptstyle\log}}{{=}}\frac{\alpha+\gamma}{2}\stackrel{{\scriptstyle\log}}{{=}}\frac{1}{2}C(a).\end{array}$
$\begin{array}[]{rcl}C(b|z)&\stackrel{{\scriptstyle\log}}{{=}}&C(b|pq)\stackrel{{\scriptstyle\log}}{{=}}C(b,p,q)-C(p,q)\\ &\stackrel{{\scriptstyle\log}}{{=}}&C(b,p)-C(p,q)\stackrel{{\scriptstyle\log}}{{=}}C(b)+C(p)-C(p)-C(q)\\ &\stackrel{{\scriptstyle\log}}{{=}}&C(b)-C(q)\stackrel{{\scriptstyle\log}}{{=}}\beta+\gamma-\frac{\beta+\gamma}{2}\stackrel{{\scriptstyle\log}}{{=}}\frac{\beta+\gamma}{2}\stackrel{{\scriptstyle\log}}{{=}}\frac{1}{2}C(b).\end{array}$
@@
$I(p:y)\stackrel{{\scriptstyle\log}}{{=}}C(p)-C(a|b)\stackrel{{\scriptstyle\log}}{{=}}\frac{\alpha+\gamma}{2}-\alpha\stackrel{{\scriptstyle\log}}{{=}}\frac{\gamma-\alpha}{2}.$
@@
  • $C(q)\stackrel{{\scriptstyle\log}}{{=}}\frac{\alpha+\beta}{2}$,
  • $C(q|b)\stackrel{{\scriptstyle\log}}{{=}}0$ (and, therefore, $I(q:b)\stackrel{{\scriptstyle\log}}{{=}}C(q)\stackrel{{\scriptstyle\log}}{{=}}\frac{\alpha+\beta}{2}$),
  • $C(q|a)\stackrel{{\scriptstyle\log}}{{=}}C(q)$ (thus, $I(q:p)\stackrel{{\scriptstyle\log}}{{\leq}}I(q:a)\stackrel{{\scriptstyle\log}}{{=}}0$).
Similarly to Case 1 we let $z=\langle p,q\rangle$ and obtain the necessary conditional complexities: the computation
$\begin{array}[]{rcl}C(a|z)\stackrel{{\scriptstyle\log}}{{=}}\frac{\alpha+\gamma}{2}\stackrel{{\scriptstyle\log}}{{=}}\frac{1}{2}C(a).\end{array}$
@@
$\begin{array}[]{rcl}C(b|z)&\stackrel{{\scriptstyle\log}}{{=}}&C(b|pq)\stackrel{{\scriptstyle\log}}{{=}}C(b,p,q)-C(p,q)\\ &\stackrel{{\scriptstyle\log}}{{=}}&C(b,p)-C(p,q)\stackrel{{\scriptstyle\log}}{{=}}C(b|p)+C(p)-C(p)-C(q)\\ &\stackrel{{\scriptstyle\log}}{{=}}&C(b)-I(p:b)-C(q)\stackrel{{\scriptstyle\log}}{{=}}\beta+\gamma-\frac{\gamma-\alpha}{2}-\frac{\alpha+\beta}{2}\stackrel{{\scriptstyle\log}}{{=}}\frac{\beta+\gamma}{2}\stackrel{{\scriptstyle\log}}{{=}}\frac{1}{2}C(b).\end{array}$
@@ A.2 Halving complexities with a topological argument
In this section, we presented an alternative (topological) proof for a special case of Theorem 1. This is the special case to which we repeatedly return as a running example throughout the paper: it involves a pair $(x,y)$ with the complexity profile
$C(x)\approx C(y)\approx 2n\quad\text{and}\quad I(x:y)\approx n.$
The argument outlined below can be combined with Muchnik’s hashing method from the previous section and extended to a significantly more general case; see [18].
Proposition 4.
@@
Then there exists a string $z$ such that $C(x|z)\stackrel{{\scriptstyle\log}}{{=}}n\text{ and }C(y|z)\stackrel{{\scriptstyle\log}}{{=}}n.$
@@
(The argument below is a simplified version of the more general proof presented in [18, Theorem 4].) We will show that the required $z$ can be constructed as a pair $a=\langle x^{\prime},y^{\prime}\rangle$, where $x^{\prime}$ and $y^{\prime}$ are prefixes of $x$ and $y$ respectively. We are going to prove that such prefixes of $x$ and $y$ can be chosen properly, although we cannot specify their lengths explicitly
@@
First of, we observe that the points $(0,0)$, $(|x|,0)$, $(0,|y|)$, $(|x|,|y|)$ are mapped to the ($o(n)$-neighborhoods of) points
@@
on the right [^3] For example, $(\alpha,\beta)=(|x|,0)$ means that $z$ is a pair consisting of the entire $x$ and the empty prefix of $y$; conditional on this $z$, complexity of $x$ vanishes, and conditional complexity of $y$ shrinks to $C(y|x)=n\pm o(n)$. The other three cases can be settled in a similar way..
When we go along the vertical line $(|x|,0)$$(|x|,|y|)$, the string $z$ consists of the entire $x$ and a gradually growing prefix of $y$. Accordingly, the value of $C(x|z)$ remains equal to $O(1)$, and $C(y|z)$ decreases from $C(y|z)\pm O(\log n)$ to 00. That is, the image of $(\alpha,\beta)$ (taken from the left picture) runs through the segment $(0,n)$$(0,0)$ (on the right picture). Similarly, when $(\alpha,\beta)$ run the segment $(0,|y|)$$(|x|,|y|)$ on the left, the images run through the segment $(n,0)$$(0,0)$ on the right, at least within $o(n)$ precision.
Now let $(\alpha,\beta)$ runs through the segment $(0,0)$$(|x|,0)$. We know that the image of $(\alpha,\beta)$ (i.e., the corresponding values of $(C(x|z),C(y|z))$) gradually changes from $(2n,2n)$ to $(0,n)$. However, we cannot control exactly the behavior of $(C(x|z),C(y|z))$ on this interval: the image of $(0,0)$$(|x|,0)$ does not need to be a straight line; we only know that it is slowly changing line that descends from $(2n,2n)$ to $(0,n)$ and remains all the way above the horizontal line $C(y|z)=n-o(n)$ (indeed, the complexity value $C(y|x^{\prime})$ cannot be be much smaller than $C(y|x)=n-o(n)$).
Similarly, when $(\alpha,\beta)$ runs through the segment $(0,0)$$(0,|y|)$, the pair of values $(C(x|z),C(y|z))$ goes from $(2n,2n)$ to $(n,0)$. Thus, the border line of the rectangle on the left in Fig. 3 is mapped to the bold line on the right in the same figure. Observe that the point $(n,n)$ is inside (at least within $o(n)$ precision) the loop shown in bold on the right in Fig. 3. It remains to show that some point from the domain of the mapping (13) is mapped to the $O(1)$-neighborhood of $(n,n)$. In what follows we do it in three steps.